# Discovering and Interpreting Shared Components for Sequence Continuation Tasks in a Large Language Model

### 13 Jan 2024 v2

@@ 1 Introduction
Figure 1: Simplified circuit show important components for the Increasing numerals (red) and Increasing Number Words (blue) tasks merged into one diagram. The purple portions denote a shared, entangled sub-circuit across both tasks. For demonstration simplicity, components exclusive to each tasks’ circuit are not shown. In §5.2.1, "Number Detection" Head 4.4 is generalized as an "Adjacent Member Detection" Head.
@@ 2 Background and Related Work
Transformer Models. We analyze LLM transformer-based models with a vocabulary size $V$. The model takes an input sequence $(\text{{x}}_{1},\ldots,\text{{x}}_{p})$ where each $\text{{x}}_{i}\in\{1,\ldots,V\}$. Tokens are mapped to $\text{{d}}_{e}$-dimensional embeddings by selecting the $\text{{x}}_{i}$-th column of $E\in\mathbb{R}^{d_{e}\times V}$ (Vaswani et al., 2017).
Attention Head. A transformer model consists of blocks of attention heads, which each consists of two matrices: the QK matrix that outputs the attention pattern $A_{i,j}\in\mathbb{R}^{N\times N}$, and the OV matrix that outputs to the residual stream. The output of an attention layer is the sum of attention heads $h_{i,j}$. We use the notation L.H for attention heads, where L is a layer index and H is a head index in layer L.
Multi-Layer Perceptron. Each attention layer output is passed to a Multi-Layer Perceptron (MLP). The MLPs in transformers are generally made of two linear layers with a ReLU activation function in between.
@@
Interpretability of Sequential Tasks. Successor heads were found by Gould et al. (2023) to increment tokens with a natural ordering. Hanna et al. (2023) found circuits for "greater-than" sequence tasks (for example completing the sentence, “The war lasted from the year 1732 to the year 17”, with any valid two-digit end years > 32). Greater-than tasks allow any year greater than a value to be valid, which differs from our sequence completion tasks that only have one valid answer. The authors noted that "similar tasks had similar… circuits", but compared across number tasks, and not across non-number tasks such as months. In our work, we study tasks that are more dissimilar.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. When we run attention pattern analysis on sequences comprised solely of sequence member tokens such as “1 2 3 4”, there are no other ‘non-sequence member’ words to compare to, so it is hard to tell what ‘type’ of token each head is attending to. Thus, we measure what types of tokens the heads attend to by using prompts that contained these sequences within other types of tokens, such as "Table lost in March. Lamp lost in April."
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a "next sequence" score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix F.
@@ 4 Discovering Circuit Connectivity
For each task, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February. Anthony was born in March. Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task; thus, there are 4608 total samples. More discussion about datasets is found in Appendix A.
@@ 4.1 Shared Sub-Circuits for Similar Sequence Continuation Tasks
We discover shared sub-circuits across the three sequence continuation tasks. Figure 2 combines all three circuits into one graph  [^5] Due to the (qkv) circuit’s large display size, we show the circuits with (qkv) decomposition in Appendix C.. These circuits were found using a performance threshold of $T_{n}=T_{e}=80$%. There is a sub-circuit found across the circuits for all three tasks, which includes heads 4.4, 7.11 and 9.1, which we show to be important in Table 2. As seen in both Figure 2 and in Table 4 in Appendix C, in which only head 5.0 of the Numerals circuit is not part of the Number Words circuit, the Numerals circuit is nearly a subset of the Number Words circuit. This suggests that the Number Words circuit uses the Numerals circuit as a sub-circuit, but requires additional components to make accurate predictions.
@@
MLP Connectivity. For all tasks, we find that several MLP ablations cause a >20% performance drop. In particular, MLP 9 causes a substantial drop of more than 90%. These results are found in Appendix D.
@@ 5.2.1 Sequence Member Detection Heads
Figure 3: Attention Patterns for (a) Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. The Months are in sequential order, but the Numerals are not. We observe for head 1.5, similar types attend to similar types. However, for head 4.4, Months attend to Months, but that Numerals do not attend to Numerals.
We discover a "similar member" detection head 1.5, and a "sequence member detection" 4.4. Attention pattern analysis reveals that these heads detect how sequence members (as queries) attend to sequence members (as keys) of the same type, such as numerals. To determine if this detection only occurs if the sequence members are in sequential order, or if this occurs even if they are not, we input prompts with Numerals in random order but with Months in sequential order. In Figure 3, we observe, for head 1.5, similar types attend to similar types. However, for head 4.4, Months attend to Months, but Numerals do not attend to Numerals, as the Numerals are not in sequential order. Therefore, in general, both heads 1.5 and 4.4 appear to detect similar token types that belong to an ordinal sequence such Numerals or Months, but head 4.4 acts even more specifically as an adjacent sequence member detection head. More discussion about these attention patterns are in Appendix G.
@@ 5.2.2 Last Token Sequence Detection Head
Figure 4: Head 7.11 Attention Pattern for Numerals. At the last token, head 7.11 has attends more to later numbers. The previous token offset pattern is in dark green.
Figure 5: The Head 9.1 Attention Pattern shows it pays strong attention to only the most recent number.
@@ 5.3 MLP Functionality
For most samples of all sequence types, logit lens reveals that the model does not predict the correct answer before MLP 9. However, after the information is processed through MLP 9, the model outputs the next sequence member. These findings suggests that MLP 9 is largely responsible for finding the next sequence member, even more so than Head 9.1, which may just be boosting this information and/or acting as a backup component. Logit lens results can be found in Appendix §D.
@@ 6 Conclusion
Understanding the inner workings of neural networks is essential for fostering alignment and safety. In this paper, we identify that across similar sequence continuation tasks, there exist shared sub-circuits that exhibit similar functionality. Specifically, we find sequence token detection heads and components associated with next sequence outputs. The aim of this work to advance our understanding of how transformers discover and leverage shared computational structures across similar tasks. By locating and comparing these circuits, we hope to gain insight into both the inductive biases that allow efficient generalization in these models and their semantic representations of abstract concepts, which may provide evidence for hierarchical associations. In future work, we plan to investigate how shared circuits affects model editing.
Limitations
@@
Dataset Size. As there are only twelve months, the number of possible continuing sequences was limited. Additionally, even if months were not used, GPT-2 also has limited prediction ability for number word sequences, as detailed in Appendix A. However, our dataset size for the months continuation task fully captures all of the months; in contrast, a small dataset size brings more issues when it does not capture all of the true distribution.
@@
Acknowledgements
We are grateful to Torr Vision Group (TVG) and Apart Lab members for feedback on the previous version, specially, Jishnu Mukhoti, Luke Marks, Clement Neo, Michelle Lo and Ashkan Khakzar.
Appendix A Dataset Details
@@
Appendix B Computational Resources and Packages
@@
Appendix C Individual Circuit Results
@@
Appendix D MLP Analysis Details
@@ 11 September, August, October
Appendix E Iterative Method Details
@@
Appendix F Functionality Method Details
@@
Appendix G Attention Pattern Extended Results
Sequence Member Detection Heads Details. We discovered a "similar member" detection head, 1.5, and a "sequence member detection", 4.4, both shown in Figure 9, where numerals attend to previous numerals, and in Figure 10, where number words attend to previous number words. Furthermore, in Figure 11, we use prompts consisting of names, same tokens ("is") and periods in the format of "<name> is <number>" (such as "Adam is 1.") to discern whether these heads are "similarity detection" heads in general, or are more specific to detecting sequence members such as numbers. This analysis shows that not all token types attend to their similar types; for instance, names do not attend to names. We also do not observe every token attending to a previous position k tokens back (where k is an integer), so we do not conclude that these heads also act as previous token heads. Additionally, Figure 12 shows that when both Numerals and Months are in sequence order, the heads attend to both Numerals and Months.
Figure 9: Attention Patterns for Increasing Digits of (a) Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. We observe that digits attend to digits.
Figure 10: Attention Patterns for Increasing Number Words of (a) Head 1.5 and (b) Head 4.4. We observe that number words attend to number words. We also observe that the attention scores here are less than they are for digits, suggesting that head 4.4 is more important for digit detection, which is consistent with its importance for the digits task over the number words task as shown in Table 2.
Figure 11: Attention Patterns for (a) Head 1.5 and (b) Head 4.4. We observe that digits attend to digits, but they are not considered general "similarity detection heads" as non-number token types do not attend to their similar or same token types.
@@ Appendix G Attention Pattern Extended Results
Appendix H Months Circuit Keeping MLP 11
@@
Appendix I Circuit Entanglement Definitions
@@
Appendix J Broader Impact Statement

### 22 Feb 2024 v3

# Interpreting Shared Circuits for Ordered Sequence Prediction in a Large Language Model

@@ 1 Introduction
Figure 1: A simplified circuit showing important components for the Numerals (red) and Number Words (blue) tasks merged into one diagram. The purple portions denote a shared, entangled sub-circuit across both tasks.
@@ 2 Background and Related Work
Transformer Models. We analyze LLM transformer-based models that take in an input sequences $(\text{{x}}_{1},\ldots,\text{{x}}_{L})$ of length $L$ Tokens are mapped to d-dimensional embeddings (Vaswani et al., 2017). An embedding is a function $\Psi:\mathcal{D}\rightarrow\mathbb{R}^{d}$, mapping token elements from a vocabulary domain $\mathcal{V}$ to vectors in $\mathbb{R}^{d}$.
Attention Head. A transformer model consists of blocks of attention heads, which each consists of two matrices: the QK matrix that outputs the attention pattern $A_{i,j}$, and the OV matrix that outputs to the residual stream. The output of an attention layer is the sum of attention heads $h_{i,j}$. We use the notation L.H for attention heads, where L is a layer index and H is a head index in layer L.
An attention head is parameterized by key, query and value matrices $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$. We denote the key, query and value outputs as $k_{i}=W_{k}\mathbf{x}_{i}$, $q_{i}=W_{q}\mathbf{x}_{i}$, $v_{i}=W_{v}\mathbf{x}_{i}$, and denote $K_{i}=[k_{1},\ldots,k_{i}]\in\mathbb{R}^{d\times i}$ and $V_{i}=[v_{1},\ldots,v_{i}]\in\mathbb{R}^{d\times i}$. The output of the head at token $i$ is denoted by $\mathbf{o}_{i}\in\mathbb{R}^{d}$, where $\mathbf{o}_{i}=V_{i}\cdot\text{softmax}(K_{i}\cdot q_{i})$.
Multi-Layer Perceptron. Each attention layer output is passed to a Multi-Layer Perceptron (MLP). The MLPs in transformers are generally made of two linear layers, parameterized by weight matrices $W_{1}$ and $W_{2}$, with a ReLU activation function in between.
Transformer-Block. An attention-block combines information from multiple attention heads, with each head attending to different parts of the input sequence. In a transformer model, $l$ attention heads are applied in parallel, forming an attention-block. Subsequently, a transformer-block processes the output of the attention-block, typically followed by an MLP. The MLP operates on the concatenated output of the $l$ attention heads, allowing the model to capture complex patterns in the data.
@@
Interpretability of Sequential Tasks. In concurrent work, successor heads were found by Gould et al. (2023) to increment tokens with a natural ordering. However, the authors do not look for circuits for sequences with more than one member, and only focus on components; in this paper, we also study shared interactions between components. Hanna et al. (2023) found circuits for "greater-than" sequence tasks (for example completing the sentence, “The war lasted from the year 1732 to the year 17”, with any valid two-digit end years > 32). Greater-than tasks allow any year greater than a value to be valid, which differs from our sequence completion tasks that only have one valid answer. The authors noted that "similar tasks had similar… circuits", but compared across number tasks, and not across non-number tasks such as months. In our work, we study tasks that are more dissimilar.
@@ 3.2 Functionality Discovery Methods
Attention Pattern Analysis. We analyze the $\mathbf{QK}$ matrix of attention heads to track information movement from keys to queries. We take the mean of datasets samples to calculate the attention scores of the heatmaps in §5.2 and §H, and display only one sample on the axes for demonstration.
Component Output Scores. We analyze head outputs by examining the values written to the residual stream via the heads’ output matrices ($\mathbf{OV}$), allowing us to see what information is being passed by each head along in the circuit. These values are measured by component output scores; we utilize a "next sequence" score that measures how well a head, given sequence token $I$, outputs token $I+1$. The details of this method are described in Appendix G.
@@ 4 Discovering Circuit Connectivity
For each task, we generate samples by placing our sequence members among non-sequence tokens. For instance, one sample may be ’Kyle was born in February. Anthony was born in March. Grant was born in April. Madison was born in’. Placing sequence members amongst non-sequence tokens allows us to evaluate the circuit representation of the shared sub-task of how the model selects sequence members from non-sequence members. We generate a total of 1536 samples per task; thus, there are 4608 total samples. More discussion about datasets is found in Appendix B.
@@ 4.1 Shared Sub-Circuits for Similar Sequence Continuation Tasks
We discover shared sub-circuits across the three sequence continuation tasks. Figure 6 in Appendix A combines all three circuits into one graph  [^5] Due to the (qkv) circuit’s large display size, we show the circuits with (qkv) decomposition in Appendix D.. These circuits were found using a performance threshold of $T_{n}=T_{e}=80$%. There is a sub-circuit found across the circuits for all three tasks, which includes heads 4.4, 7.11 and 9.1, which we show to be important in Table 2. This is shown in Figure 2. As seen in both Figure 6 and in Table 4 in Appendix D, in which only head 5.0 of the Numerals circuit is not part of the Number Words circuit, the Numerals circuit is nearly a subset of the Number Words circuit. This suggests that the Number Words circuit uses the Numerals circuit as a sub-circuit, but requires additional components to make accurate predictions.
Figure 2: The important sub-circuit components for all three tasks. Resid_post denotes the residual stream state right before the linear unembedding to logits. The full circuits are shown in Appendix A.
@@
MLP Connectivity. For all tasks, we find that several MLP ablations cause a >20% performance drop. In particular, MLP 9 causes a substantial drop of more than 90%. These results are found in Appendix E.
@@ 5.2.1 Sequence Member Detection Heads
Figure 3: Attention Patterns for (a) Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. The Months are in sequential order, but the Numerals are not. We observe for head 1.5, similar types attend to similar types. However, for head 4.4, Months attend to Months, but that Numerals do not attend to Numerals. Note we take the mean of from dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration.
We discover a "similar member" detection head 1.5, and a "sequence member detection" 4.4. Attention pattern analysis reveals that these heads detect how sequence members (as queries) attend to sequence members (as keys) of the same type, such as numerals. To determine if this detection only occurs if the sequence members are in sequential order, or if this occurs even if they are not, we input prompts with Numerals in random order but with Months in sequential order. In Figure 3, we observe, for head 1.5, similar types attend to similar types. However, for head 4.4, Months attend to Months, but Numerals do not attend to Numerals, as the Numerals are not in sequential order. Therefore, in general, both heads 1.5 and 4.4 appear to detect similar token types that belong to an ordinal sequence such Numerals or Months, but head 4.4 acts even more specifically as an adjacent sequence member detection head. More discussion about these attention patterns are in Appendix H.
@@ 5.2.2 Last Token Sequence Detection Head
Figure 4: Attention Pattern for Head 7.11. At the last token, head 7.11 to more recent sequence members than earlier sequence members. We also recognize an offset pattern on the diagonals of this heatmap, indicating that it also functions as a previous token head. Note we take the mean of from the Numerals dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration.
Figure 5: The Attention Pattern of Head 9.1 shows the component pays strong attention to only the most recent number. Note we take the mean of from the Numerals dataset samples to calculate the attention scores, but display only one sample on the axes for demonstration.
@@ 5.3 MLP Functionality
For most samples of all sequence types, logit lens reveals that the model does not predict the correct answer before MLP 9. However, after the information is processed through MLP 9, the model outputs the next sequence member. These findings suggests that MLP 9 is largely responsible for finding the next sequence member, even more so than Head 9.1, which may just be boosting this information and/or acting as a backup component. Logit lens results can be found in Appendix §E.
@@ 6 Conclusion
Understanding the inner workings of neural networks is essential for fostering alignment and safety. In this paper, we identify that across similar sequence continuation tasks, there exist shared sub-circuits that exhibit similar functionality. Specifically, we find sequence token detection heads and components associated with next sequence outputs. The aim of this work is to advance our understanding of how transformers leverage shared computational structures across similar tasks. By locating and comparing these circuits, we hope to gain insight into semantic representations of abstract concepts, which may provide evidence for hierarchical associations.
7 Limitations
@@
Dataset Size. As there are only twelve months, the number of possible continuing sequences was limited. Additionally, even if months were not used, GPT-2 also has limited prediction ability for number word sequences, as detailed in Appendix B. However, our dataset size for the months continuation task fully captures all of the months; in contrast, a small dataset size brings more issues when it does not capture all of the true distribution.
@@
Appendix A All Three Circuits Diagram
In Figure 6, we show the three circuits described in §4 found via iterative node and edge ablation. Due to its large size with many edges, we place it in the Appendix rather than in the main paper.
Appendix B Dataset Details
@@
Appendix C Computational Resources and Packages
@@
Appendix D Individual Circuit Results
@@
Appendix E MLP Analysis Details
@@ 11 September, August, October
Appendix F Iterative Method Details
@@
Appendix G Functionality Method Details
@@
Appendix H Attention Pattern Extended Results
When we run attention pattern analysis on sequences comprised solely of sequence member tokens such as “1 2 3 4”, there are no other ‘non-sequence member’ words to compare to, so it is hard to tell what ‘type’ of token each head is attending to. Thus, we measure what types of tokens the heads attend to by using prompts that contained these sequences within other types of tokens, such as "Table lost in March. Lamp lost in April."
Sequence Member Detection Heads Details. We discovered a "similar member" detection head, 1.5, and a "sequence member detection", 4.4, both shown in Figure 10, where numerals attend to previous numerals, and in Figure 11, where number words attend to previous number words. To discern whether these heads are "similarity detection" heads in general, or are more specific to detecting sequence members such as numbers, these figures show that not all token types attend to their similar types; for instance, names do not attend to names. We also do not observe every token attending to a previous position k tokens back (where k is an integer), so we do not conclude that these heads also act as previous token heads. Additionally, Figure 12 shows that when both Numerals and Months are in sequence order, the heads attend to both Numerals and Months.
Figure 10: Attention Patterns for Numerals of (a) Head 1.5 and (b) Head 4.4. Lighter colors mean higher attention values. For each of these detection patterns, the query is shown in green, and the key is shown in blue. We observe that numerals attend to numerals, but they are not considered general "similarity detection heads" as non-number token types do not attend to their similar or same token types.
Figure 11: Attention Patterns for Number Words of (a) Head 1.5 and (b) Head 4.4. We observe that number words attend to number words (for head 4.4, these scores are in dark blue). We also observe that the attention scores here are less than they are for digits, suggesting that head 4.4 is more important for digit detection, which is consistent with its importance for the digits task over the number words task as shown in Table 2.
@@ Appendix H Attention Pattern Extended Results
Appendix I Months Circuit Keeping MLP 11
@@
Appendix J Circuit Entanglement Definitions
@@
Appendix K Broader Impact Statement